# MambaXCTrack: Mamba-based Tracker with SSM Cross-correlation and Motion Prompt for Ultrasound Needle Tracking

### v1

@@ IV CONCLUSIONS
In this paper, a US needle tracker MambaXCTrack with SSMX-Corr, CIS, and low-level motion descriptor has been introduced to address needle tracking problems under challenging US imaging environments with noise, artifacts, and intermittent tip visibility. It is the first Mamba-based tracker for US needle tracking. Experiments and ablation studies on a dataset of motorized insertions in both phantom and tissue show its effectiveness. MambaXCT outperforms the other SOTA methods in almost all metrics in all evaluations. One of the limitations is the efficiency, where it requires a high-end GPU to achieve satisfactory inferencing speed. In our future work, improvements will be made for a more efficient X-Corr structure based on SSM. More experiments on needle tracking during manual needle insertion will also be conducted to validate the tracking robustness.

### v2

@@ IV CONCLUSIONS
In this paper, a US needle tracker MambaXCTrack with SSMX-Corr, CIS, and a low-level motion descriptor has been introduced for US needle tracking under challenging US imaging environments with noise, artifacts, and intermittent tip visibility. To the best of our knowledge, it is the first Mamba-based tracker for US needle tracking. Experiments and ablation studies show its effectiveness. However, MambaXCTrack requires a high-end GPU to achieve real-time inference, hindering deployment on low-end hardware. In future work, improvements will be made for better efficiency, including network pruning, model distillation, parallel optimization, etc. More experiments on manual needle insertion will also be conducted. The code implementation of the proposed method will be made publicly available in future work once ongoing developments are completed.